Use the bond term's below to answer the question

Maturity 7 years

Coupon Rate 4%

Face value $1,000

Annual Coupons

YTM 5%

Assuming the YTM remains constant throughout the bond's life, what is the bond's current yield between periods 2 and 3 ?

4.18%
4.10%
4.31%
3.98%
Answer #1

Coupon = 0.04 * 1000 = 40

Price = Coupon * [1- 1 / (1 + r)n] / r + FV / (1 + r)n

Price = 40 * [1 - 1 / (1 + 0.05)5] / 0.05 + 1000 / (1 + 0.05)5

Price = 40 * 4.329477 + 783.526166

Price = 956.7052

Current yield = (Annual coupon / price) * 100

Current yield = (40 / 956.7052) * 100

Current yield = 4.18%